# Cluster Catch Digraphs with the Nearest Neighbor Distance

### v1

@@ Abstract
We introduce a new method for clustering based on Cluster Catch Digraphs (CCDs). The new method addresses the limitations of RK-CCDs by employing a new variant of spatial randomness test that employs the nearest neighbor distance (NND) instead of the Ripley’s $K$ function used by RK-CCDs. We conduct a comprehensive Monte Carlo analysis to assess the performance of our method, considering factors such as dimensionality, data set size, number of clusters, cluster volumes, and inter-cluster distance. Our method is particularly effective for high-dimensional data sets, comparable to or outperforming KS-CCDs and RK-CCDs that rely on a KS-type statistic or the Ripley’s $K$ function. We also evaluate our methods using real and complex data sets, comparing them to well-known clustering methods. Again, our methods exhibit competitive performance, producing high-quality clusters with desirable properties.
@@
Keywords: Graph-based clustering, Cluster catch digraphs, High-dimensional data, The nearest neighbor distance, Spatial randomness test.
1 Introduction
@@
Clustering is a fundamental statistical and data analysis method, partitioning observations into groups based on similarity or proximity. Clustering is considered an unsupervised version of classification, and the primary goal is to understand the inherent structure of data, which plays a crucial role in machine learning and pattern recognition. Clustering faces challenges like determining the optimal number of clusters and handling high-dimensional data. Driver and Kroeber first introduced cluster analysis in 1932 [15], applying it to anthropology. Early clustering methods focused on $k$-means and hierarchical methods, which are the fundamental of cluster analysis [54, 35]. Since then, clustering methods have evolved substantially and are widely used across different disciplines. Common examples include partitioning customers into distinct groups for better market strategy [60], grouping genes that have a similar physiological expression in Biology [18], enhancing image processing by dividing and re-grouping [12], and predicting human behavior based on their communities and social ties [25]. Despite significant advancements, clustering remains challenging, such as determining the optimal number of clusters [39], scalability to large data sets [67], robustness to outliers and noise [51], and handling high-dimensional data [6].
We propose clustering methods based on Cluster Catch Digraphs (CCDs). CCDs were first introduced by DeVinney [14] and Marchette [38], developed from a similar classification digraph called Class Cover Catch Digraphs (CCCDs). Later, Manukyan and Ceyhan [36] modified and improved this approach further, developing two variants using a Kolmogorov-Smirnov (KS) based statistic and Ripley’s $K$ function, calling them KS-CCDs and RK-CCDs, respectively, RK-CCDs and KS-CCDs work similarly in clustering, and RK-CCDs are almost parameter-free, making them especially appealing. However, our experimental analysis shows that RK-CCDs may not be suitable for moderate to high dimensionality due to the limitation of the Spatial Randomness Monte Carlo Test (SR-MCT) based on Ripley’s $K$ function [53]. Thus, we introduce another CCD-based approach that uses Nearest Neighbors Distance (NND) instead of Ripley’s $K$ function to test underlying point-process patterns, and the resulting methods are called UN-CCDs [53]. Like RK-CCDs, UN-CCDs are almost parameter-free. We introduce the U-MCCD and SU-MCCD methods for outlier detection[53], which are based on RK-CCDs and UN-CCDs, respectively. We have shown that the SU-MCCD method performs comparably to the U-MCCD method when the number of dimensions $d\leq 5$, and outperforms it substantially when $d\geq 10$. Then, we propose two “flexible” variations of U-MCCDs and SU-MCCDs, called UN-MCCDs and SUN-MCCDs, respectively, aiming at outlier detection on the data sets with arbitrary-shaped clusters and varying intensities.
2 Literature and Background Review
So far, cluster-based methods have been classified into several subgroups, known as partitional, hierarchical, density-based, grid-based, fuzzy-based, model-based, and distribution-based clustering methods.
2.1 Partitioning-Based Clustering Methods
Partitioning or distance-based clustering methods follow an iterative process [57], dividing a data set into single-level partitions, which differ from the hierarchical structure obtained by other clustering techniques [62]. These methods typically start with a predefined number of clusters, often represented by their centers, obtained through a simple method like random selection. The partitions are then iteratively refined until a specific criterion, typically the sum of squares of within-cluster (or inter-cluster) distances, is minimized (maximized). Partition-based clustering methods divide a data set into meaningful partitions for subsequent analysis [62]. While efficient and simple, they are sensitive to outliers and require the number of clusters as an input parameter [62]. Commonly known partition-based clustering methods include $K$-Means [24], MacQueen [35], Partitioning Around Medoids (PAM) [33], Clustering LARge Applications (CLARA) [33], and Clustering Large Applications based on RANdomized Search (CLARANS) [41].
2.2 Hierarchical Clustering Methods
The performance of partitioning-based clustering methods depends on the chosen number of clusters $k$. Determining the optimal value for $k$ can be difficult when there is limited knowledge about the data set. Fortunately, hierarchical clustering methods offer a solution to this issue. Typically, these methods construct a hierarchical tree-like structure known as dendrogram, which allows for partitioning the whole data set at varying levels of granularity. Hierarchical clustering methods can be classified into two categories: agglomerative and divisive clustering [65]. Agglomerative clustering begins by treating each point as a cluster, merging the two closest clusters iteratively until a pre-defined stopping criterion is satisfied. Divisive clustering operates oppositely, treating the entire data set as a single cluster and partitioning iteratively.
Hierarchical clustering is versatile and can be applied to various types of data. The dendrogram provides a visual summary of the clustering process, increasing interpretability. However, hierarchical clustering can be computationally expensive, especially for large data sets [45]. Commonly known hierarchical clustering methods includes minimum variance method [59], Minimal Spanning Tree [63], Clustering Using Representatives (CURE) [26], CHAMELEON [32], and RObust Clustering using links (ROCK) [27].
2.3 Density-Based Clustering Methods
The density-based clustering methods work similarly to the human eye, which identifies regions with high concentrations of data points as clusters. Like the hierarchical clustering methods, these methods do not require specifying the number of clusters. They can capture clusters of arbitrary shapes and sizes, exhibiting robustness to outliers. However, their performance depends on the input density parameters. Moreover, they may have difficulty identifying clusters under complex and high-dimensional data sets [20]. Recent density-based clustering methods include Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [19], Ordering Points To Identify the Clustering Structure (OPTICS) [2], Distribution Based Clustering of LArge Spatial Databases (DBCLASD), and DENsity-based CLUstEring (DENCLUE) [28].
2.4 Grid-Based Clustering Methods
Grid-based clustering methods operate by partitioning the data space into a finite number of cells, creating a grid structure. Clusters are then constructed based on these cells. Generally, the computational complexity depends on the granularity of the grid rather than the data size. Therefore, they can be efficient with large data sets and are suitable for arbitrarily shaped clusters thanks to their flexibility. However, the quality of clustering depends on the granularity. Moreover, they have difficulty capturing clusters with irregular boundaries [62]. Some well-known grid-based clustering methods include STING [58], WaveCluster [52], DClust [66], and Clustering In QUEst (CLIQUE) [1].
2.5 Model-Based Clustering Methods
Model-based (or distribution-based) clustering methods operate under the assumption that data is generated from a mixture of distributions, with each component representing a cluster. These methods aim to recover the models and parameters that best fit the observed data, constructing clusters based on the estimated models [45]. The major advantage of them is their flexibility to the cluster shapes and sizes, without requiring a predetermined number of clusters. Moreover, they allow “soft” clustering, where a data point can belong to multiple clusters. However, model-based clustering methods can be computationally expensive, and the clustering quality highly depends on the parameters of the chosen models [62]. Some examples include Gaussian Mixture Models (GMMs) [46], Latent Dirichlet Allocation (LDA) [7], Hidden Markov Models (HMM) [44], Expectation Maximization (EM) [13], COBWEB [21], pdfCluster [4], and SOM [56].
2.6 Fuzzy Clustering
Fuzzy clustering methods, also known as “soft” clustering, allow data points to belong to two or more clusters simultaneously, each with an associated degree of membership [20]. This allows clusters to overlap and makes fuzzy clustering suitable for data sets with ambiguous and poorly separated cluster boundaries [33]. Additionally, these methods offer insights into the underlying data structure. However, the iterative optimization process on membership degrees can be computationally expensive [20].
We enumerate some fuzzy clustering methods: Fuzzy C-means (FCM) is another variant of $K$-Means and allows soft clustering [17] . However, it is sensitive to initial partition, not robust to outliers and noise, and can get trapped in a local optimum [50]. Possibilistic-fuzzy C-means (PFCM) combines possibilistic C-means (PCM) and FCM [42, 34]. Therefore, it is more robust to noise and outliers. Kernelized Fuzzy C-means (KFCM) is another variant of FCM [64], incorporating the kernel method to handle clusters with non-linear boundaries.
@@ 3.1 Cluster Catch Digraphs
Our clustering methods are based on Cluster Catch Digraphs (CCDs), where vertices represent the points of a given data set $\mathcal{X}$, and arcs are determined by the spherical ball (covering ball) centered at each point. Class Cover Catch Digraphs (CCCDs) [43] are the prototypes of CCDs, originally designed for supervised classification. CCCDs aim to distinguish a target class from a non-target class by covering the target class points with a minimum number of open balls while excluding the non-target points [38]. A CCCD is constructed by creating covering balls around each point, forming arcs between points when one lies within the covering ball of another. Two variants, pure-CCCDs (P-CCCDs) and random walk-CCCDs (RW-CCCDs), differ in how the radius of each covering ball is determined [36]. As in the CCCD classification, to reduce the covering complexity in CCDs and avoid overfitting, we retain only a subset of covering balls that optimally covers all (or most) points in a data set. One way is finding a minimum dominating set (MDS) of the CCD. Although finding the MDS of a digraph is an NP-Hard problem [31], an efficient approximation of MDS can be found using the Greedy Algorithm, which operates in $O(|V_{0}|^{2})$ time [9, 29]. The Greedy Algorithm finds an approximate MDS by iteratively selecting the vertex with the highest outdegree, adding it to the dominating set, and removing its closed neighborhood (in the graph) from the uncovered vertices until all vertices are covered. Manukyan and Ceyhan proposed two variants of the Greedy Algorithm for CCDs [37]: the first variant prioritizes vertices closer to cluster centers, while the second maximizes a score function $sc(v)$ at each step. We provide these methods in Algorithms 1 and 2, respectively, since they are non-standard in their greediness and also for future reference.
Input: A digraph $D=(V(D)=\mathcal{X},A(D))$ where $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}$
@@
1 Initialization: $V_{temp}\leftarrow V(D)$, $V\leftarrow V(D)$, $A\leftarrow A(D)$, $\hat{S}\leftarrow\emptyset$
2 while $V_{temp}\neq\emptyset$ do
@@ 6 $D\leftarrow D^{sub}(V_{temp})$;
Greedy Algorithm 1 (A greedy algorithm finding an approximate MDS) This greedy algorithm is adapted for CCDs. $V$ and $A$ are the initial vertex and edge sets of the digraph $D$; $D^{sub}(S)$ is the induced sub-digraph of vertex set $S$ from a digraph $D$, $\bar{N}(v)$ is the closed neighborhood of a vertex $v$ (i.e., the points covered by the covering ball centered at $v$). $V_{temp}$ represents the uncovered vertices at current iteration.
Input: A digraph $D=(V(D),A(D)).$ for a given data set $\mathcal{X}=\{x_{1},x_{2},...,x_{n}\}$
@@
1 Initialization: $V_{temp}\leftarrow V(D)$, $\hat{S}\leftarrow\emptyset$
2 while $V_{temp}\neq\emptyset$ do

### v2

@@ Abstract
We introduce a new clustering method based on Cluster Catch Digraphs (CCDs). The new method addresses the limitations of RK-CCDs on clustering by employing a new variant of spatial randomness test that employs the nearest neighbor distance (NND) instead of the Ripley’s $K$ function. Our method is particularly effective for high-dimensional datasets, comparable to or outperforming KS-CCDs and RK-CCDs (rely on a KS-type statistic or the Ripley’s $K$ function) on clustering. We conduct a comprehensive Monte Carlo analysis to assess the performance of our method, considering factors such as dimensionality, dataset size, number of clusters, cluster volumes, and inter-cluster distance. We also evaluate our methods using real and complex datasets, comparing them to well-known clustering methods. Again, our methods exhibit competitive performance, producing high-quality clusters with desirable properties.
@@
Keywords: Graph-based clustering; Cluster catch digraphs; High-dimensional data; the nearest neighbor distance; Spatial randomness test.
1 Introduction and Motivation
@@
Clustering is a fundamental Statistical method for partitioning observations into groups based on similarity or proximity. It is an unsupervised learning method that plays a crucial role in machine learning and pattern recognition. The primary goal is to group a dataset so that points in the same group (called a cluster) are similar. Clustering faces challenges like determining the optimal number of clusters and handling high-dimensional data. Driver and Kroeber first introduced cluster analysis in 1932 [driver1932quantitative], applying it to anthropology. Early clustering methods focused on $k$-means and hierarchical methods, which are the fundamentals of cluster analysis [sorensen1948method, macqueen1967classification]. Since then, clustering methods have evolved substantially and are widely used across various disciplines. Common examples include partitioning customers into distinct groups for better market strategy [wedel2000market], grouping genes that have a similar physiological expression in biology [eisen1998cluster], enhancing image processing by dividing and re-grouping [comaniciu2002mean], and predicting human behavior based on their communities and social ties [girvan2002community]. Despite significant advancements, clustering still remains challenging, due to the fact including determining the optimal number of clusters [milligan1985examination], scalability to large datasets [zhang1996birch], robustness to outliers and noise [schubert2017dbscan], and handling high-dimensional data [beyer1999nearest].
In our previous work, We have proposed a graph- and distribution-based clustering method, called UN-CCDs [shi2024outlier], aiming to overcome the limitations of its two predecessors: KS-CCDs and RK-CCDs. Specifically, RK-CCDs are ineffective for high-dimensional clustering due to Ripley’s $K$ function in MC-SRT; and KS-CCDs are not parameter-free, an intensity parameter $\delta$ is needed for a KS-type statistic. We first recap the CCD-based method below.
Cluster Catch Digraphs (CCDs) are a family of graph-based clustering methods, which were first introduced by DeVinney [devinney2003class] and Marchette [marchette2005random], adapted from a similar classification digraph called Class Cover Catch Digraphs (CCCDs). Later, Manukyan and Ceyhan [manukyan2019parameter] advanced this approach further, developing two variants using a Kolmogorov-Smirnov (KS)-type statistic and Ripley’s $K$ function, calling them KS-CCDs and RK-CCDs, respectively. RK-CCDs and KS-CCDs work similarly on clustering. While both are effective, RK-CCDs are particularly appealing for being “almost” parameter-free, unlike KS-CCDs, which require an input intensity parameter for the KS-type statistic.
However, despite their advantages, RK-CCDs are not suitable for clustering moderate to high-dimensional data, as shown by our intensive simulation study [shi2024outlier]. This gap motivates the need for a new CCD-based clustering method that is robust in higher dimensions, while remaining parameter-free.
To address this challenge, we had made the following contribution in this paper:
  • (i)
    We proposed a new clustering method called UN-CCDs (Uniformity- and Neighbor-based CCDs), which is “almost” parameter-free.
  • (ii)
    UN-CCDs address the limitations of RK-CCDs by using the nearest neighbor distance (NND) for a spatial randomness test instead of the Ripley’s $K$ function, Making this method more effective for high-dimensional datasets.
  • (iii)
    By extensive Monte Carlo analysis, We show that UN-CCDs are competitive with or superior to other well-known traditional or recent clustering methods.
  • (iv)
    The UN-CCD method can provide high clustering quality under different levels of noise.
The article is organized as follows: (1) In Section 2, we review the state-of-the-art clustering methods from different categories, plus the mathematical foundation of KS-CCDs and RK-CCDs. (2) Section 3 details the limitations of RK-CCDs first, followed by the motivation and the algorithmic formulation of UN-CCDs. (3) In Section 4 and 5, we apply UN-CCDs to real and complex datasets, demonstrating its advantages over RK-CCDs, KS-CCDs, and traditional clustering methods such as DBSCAN, $K$-means++, etc.
@@ 2 Preliminaries
2.1 Cluster Methods in Literature
In literature, cluster-based methods have been classified into several subgroups, known as partitional, hierarchical, density-based, grid-based, fuzzy-based, model-based, and distribution-based clustering methods. We provide a short summary of these method below. [wang2019survey, yin2024rapid, ezugwu2022comprehensive].
Partitioning clustering methods iteratively partition data into a predefined number ($k$) of clusters, often minimizing within-cluster distances [wang2019survey, yin2024rapid]. They are generally efficient but require $k$ as input and are sensitive to outliers [yin2024rapid]. Classic examples include $K$-Means [gareth2013introduction], $K$-Means++ [arthur2006k], MacQueen [macqueen1967classification]. PAM [kaufman2009finding] is more robust to outliers, but computationally expensive [yin2024rapid]. More recent methods, such as CLARA [kaufman2009finding] and CLARANS [ng2002clarans], address the computational cost of PAM on large datasets.
Hierarchical clustering methods construct a dendrogram with agglomerative (bottom-up) or divisive (top-down) methods, without the need to pre-define the number of clusters. Those methods offer interpretability but can be computationally costly for large datasets [ran2023comprehensive]. Classic examples include the Ward’s minimum variance method [ward1963hierarchical] and minimal spanning tree [MST], but they are computationally intensive and struggle with arbitrarily-shaped clusters [zhang2013advancements]. More recent methods, such as CURE [cure], CHAMELEON [karypis1999chameleon], and ROCK [guha2000rock] aim to overcome these limitations.
Density-based clustering methods identify clusters as high-density regions, separating them by low-density areas. They typically do not require $k$ as an input parameter, capable to find arbitrarily shaped clusters, and are robust to outliers [ester1996density]. A typical classic example is DBSCAN [ester1996density], but it’s performance is sensitive to the values of two input parameters (Eps and MinPts), and has limitations on clusters with varying intensities. Later methods were developed to solve one or two disadvantages of DBSCAN, including OPTICS [ankerst1999optics], DBCLASD [xu1998distribution], and DENCLUE [hinneburg1998efficient].
Graph-based clustering methods partition nodes (points) by leveraging a graph’s connectivity structure, aiming to identify sub-graphs for latent clusters. A key advantage of those methods is their ability to identify non-convex clusters where vector-based approaches may fail [saxena2017review]. Well-known classic examples include the Minimal Spanning Tree (MST) [MST], spectral clustering [ng2001spectral], but these methods are either sensitive to outliers or computationally intensive. More recent and scalable methods include Louvain [blondel2008fast] and Leiden [traag2019louvain], which are designed for large datasets. We also have deep learning methods like Structural Deep Clustering Network (SDCN) [bo2020structural]. This paper focuses on CCD-based methods, specifically RK-CCDs and KS-CCDs [manukyan2019parameter], which will be discussed in detail shortly. In Section 5, we will use real-life datasets to compare UN-CCDs with some commonly used graph-based methods, including MST, spectral clustering, and Louvain. We will discuss the mechanisms and setup of those methods in Section 5.
Recent work in graph clustering has focused on tackling the challenges of high-dimensional data and developing more robust methods. The “curse of dimensionality” poses a significant challenge for traditional clustering algorithms, as distances between any pairs of points become meaningless in high-dimensional spaces. UN-CCDs address this by employing a spatial randomness test with several key enhancements compared to the one employed in RK-CCDs, making it more robust to high-dimensional data. Recent research in this area also focuses on high-dimensional data and using robust statistics to identify and handle outliers that can negatively impact performance. Methods like RSC-SGM (Robust Spectral Clustering for Sub-Gaussian Mixtures) construct a Gaussian-kernel similarity graph, apply a denoising step to the affinity matrix, and use a rounding procedure to achieve accurate and automatic clustering and outlier detection [srivastava2023robust]. NA-COV (Network-Adjusted Covariates) integrates high-dimensional node covariates with network structure to improve community detection in sparse graphs [hu2024network]. BayesCov-SBM (Bayesian Community Detection with Covariates) uses a joint model called Bayesian stochastic block model that incorporates both network structure and high-dimensional node covariates for community detection [shen2025bayesian].
Graph-based clustering methods, including the CCD-based framework, Leverage the digraph’s connectivity and relative distance for clustering. A key advantage of CCD-based methods is identifying non-convex clusters [manukyan2019parameter], where vector-based approaches like $K$-Means and Hierarchical clustering may fail. Additionally, unlike density-based methods like DBSCAN that identify the “seeds” of clusters in high-density regions exceeding a density threshold, most CCDs do not require the user to define such a threshold. Instead, CCDs create a digraph where arcs are determined by a spatial randomness test for each data point, effectively identifying local clusters. CCD-based methods also differ from modularity-based graph methods, such as Louvain, which partitions a network by maximizing a measure of community structure. While Louvain and other similar methods are designed for large networks, CCDs are particularly focused on using geometric properties of the data itself to define the graph structure and clusters.
Several modern categories of clustering, including grid-based, model-based, and fuzzy-based clustering methods were developed to address specific challenges like large-scale data, statistical rigor, and ambiguous cluster boundaries.
Grid-based clustering methods partition the data space into a grid structure and perform clustering on the grid cells. Their time/space complexity depends on grid size rather than data size, making them fast for large datasets [wang1997sting]. The clustering quality relies on grid granularity, and they may struggle with irregular cluster boundaries [yin2024rapid]. Examples include STING [wang1997sting], WaveCluster [sheikholeslami2000wavecluster], DClust [zhang2005clustering], and CLIQUE [agrawal1998automatic].
Model-based clustering methods assume data originates from a mixture of probability distributions, these methods fit models to identify clusters [ran2023comprehensive]. They offer flexibility in cluster shape/size and support “soft” assignments but can be computationally intensive and are sensitive to model/parameter choices [yin2024rapid]. Examples include GMMs [rasmussen1999infinite], LDA [blei2003latent], HMM [rabiner1986introduction], COBWEB [fisher1987knowledge], pdfCluster [azzalini2013clustering], and SOM [vesanto2000clustering].
Fuzzy Clustering is known as “soft” clustering, these methods allow data points to belong to multiple clusters with varying degrees of membership [ezugwu2022comprehensive], useful for ambiguous or overlapping boundaries [kaufman2009finding]. Optimization can be computationally intensive [ezugwu2022comprehensive]. Examples include Fuzzy C-means (FCM) [dunn1973fuzzy], which is sensitive to initialization and outliers [saxena2017review]; the more robust Possibilistic-fuzzy C-means (PFCM) [pal2005possibilistic, krishnapuram1996possibilistic]; and Kernelized Fuzzy C-means (KFCM) for non-linear boundaries [zhang2004novel].
@@ 2.2 Cluster Catch Digraphs
Our clustering methods are based on Cluster Catch Digraphs (CCDs), where vertices represent the points of a given dataset $\mathrm{X}$, and arcs are determined by the spherical ball (covering ball) centered at each point. Class Cover Catch Digraphs (CCCDs) [marchette2003classification] are the prototypes of CCDs, originally designed for supervised classification. CCCDs aim to distinguish a target class from a non-target class by covering the target class points with a minimum number of balls while excluding the non-target points [marchette2005random]. A CCCD is constructed by creating covering balls around each point, forming arcs between points when one lies within the covering ball of another. Two variants, pure-CCCDs (P-CCCDs) and random walk-CCCDs (RW-CCCDs), differ in how the radius of each covering ball is determined [manukyan2016classification]. CCDs are an unsupervised adaptation of CCCDs [marchette2005random], initializing by constructing a covering ball for each point. To reduce the covering complexity in CCDs and avoid overfitting, we retain only a subset of covering balls that optimally covers all (or most) points in a dataset. One way is finding a minimum dominating set (MDS) of the CCD. Although finding the MDS of a digraph is an NP-Hard problem [karp1972reducibility], one can find an approximation of MDS efficiently with the Greedy Algorithm, which operates in $O(|V_{0}|^{2})$ time [chvatal1979greedy, hochbaum1982approximation]. The Greedy Algorithm finds an approximate MDS by iteratively selecting the vertex with the highest outdegree at each iteration and storing this point in a set, and removing its closed neighborhood (in the graph) from the uncovered vertices until all vertices are covered. When the algorithm terminates, the stored vertices are retained for the (approximate) MDS. Manukyan and Ceyhan proposed two variants of the Greedy Algorithm for CCDs [manukyan2019parameter]: the first variant prioritizes vertices closer to cluster centers, while the second maximizes a score function at each step. We provide these methods in Algorithms 1 and 2, respectively for completeness, because they are non-standard in their greediness and also for future reference.
Input: A digraph $D=(V(D)=\mathrm{X},A(D))$ where $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}$
@@
18pt Initialization: $V_{temp}\leftarrow V(D)$, $\hat{S}\leftarrow\emptyset$
25pt while $V_{temp}\neq\emptyset$ do
@@
Greedy Algorithm 1 A greedy algorithm to find an approximate MDS of CCDs (which does not update the digraph after each iteration). $V$ and $A$ are the initial vertex and edge sets of the digraph $D$, respectively. $\bar{N}(v)$ is the closed neighborhood of a vertex $v$ (i.e., the points covered by the covering ball centered at $v$, including $v$ itself). At each iteration, $V_{temp}$ represents the uncovered vertices, and $v_{temp}$ is the vertex with largest outdegree.
Input: A digraph $D=(V(D),A(D)).$ for a given dataset $\mathrm{X}=\{x_{1},x_{2},...,x_{n}\}$
@@
18pt Initialization: $V_{temp}\leftarrow V(D)$, $\hat{S}\leftarrow\emptyset$
25pt while $V_{temp}\neq\emptyset$ do